OpenStudy (venomblast):

someone please explain this: Explain in words what 60 0 1 ( ) 100 60 0 F t dt    means if () Ftis the rate, in thousands of people per year, at which a country’s population has been growing and t is time, in years.

OpenStudy (venomblast):

what is F(t). There an error while copying the problem. It should show this \[\frac{ 1 }{ 60-0 }\int\limits_{0}^{60}F(t)dt \approx 100\]

OpenStudy (baru):

do you know the graphical significance of integration?

OpenStudy (venomblast):

yes But I need to know what F(t) right?

OpenStudy (baru):

F(t) is the function that rate at which the population of some country is growing. the question is asking you to interpret the meaning of this: \[\frac{ 1 }{ 60-0 }\int\limits\limits_{0}^{60}F(t)dt \approx 100\] the question is not asking you to find the function F(t), it is only asking you to interpret meaning of the above expression.

OpenStudy (venomblast):

I know but what I am trying to say is , F(t) is irrelevant?

OpenStudy (venomblast):

I suck at interpreting

OpenStudy (baru):

for this question, all we need to know is that F(t) represents rate at which the country's population is growing. We do not need to know what the actual function is.

OpenStudy (baru):

so, do you know the graphical significance of integration?

OpenStudy (venomblast):

somewhat.

OpenStudy (baru):

suppose i'm driving a car the speed of my car is given by the function v(t) can you tell me what physical quantity we get when we do \[\int\limits v(t) dt\] ?

OpenStudy (venomblast):

a(t)?

OpenStudy (baru):

no...

OpenStudy (venomblast):

isnit it acceleration. the derivative of acceleration is velocity.

OpenStudy (baru):

the derivative of velocity is acceleration.... this is not something that needs to be memorized... it becomes clear if you understand the meaning of derivative and integral

OpenStudy (venomblast):

I do know them. I take higher math course than calculus. prepping for exam P.

OpenStudy (venomblast):

Ok you are right. I got it mixed up.

OpenStudy (baru):

ok :) so can you give this another try?\[\int\limits\limits v(t) dt\]

OpenStudy (will.h):

@baru Congrats on Mathlete title. :)

OpenStudy (baru):

haha thank you @Will.H :D

OpenStudy (venomblast):

Im not sre

OpenStudy (venomblast):

i give up. this is going nowhere

OpenStudy (mathmale):

Please note that your initial "60 0 1 ( ) 100 60 0 F t dt    " is practically unintelligible. You'd get faster and better results by drawing the integral in question, as Baru attempted to do (above).

OpenStudy (mathmale):

Here you're being asked to interpret what the given integral represents.